The length of a rectangle is 8.03 inches more than half of its width. If the perimeter is 37.48 inches
Find the demensions

Answer by addingup(3677)

2L + 2W = Perimeter
L = 1/2W + 8.03 substitute for L above:
2(1/2W + 8.03) + 2W = 37.48
W + 16.06 + 2W = 37.48
3W = 21.42
W = 7.14
L = 1/2W + 8.03 = 1/2(7.14)+8.03 = 11.6
-----------------------
Check:
2L + 2W = 37.48
2(7.14) + 2(11.6) = 37.48 Correct